The table shows the output of chairs at a factory when different numbers of workers are employed.


Number of workers 1 2 3 4 5


Number of chairs produced 6 17 27 32 30


Show where diminishing marginal returns to labour will set in.

1
Expert's answer
2021-12-01T10:14:03-0500

MP(n)=TP(n)-TP(n-1)

where;

n=number of input


TP=total product


MP=marginal product


MP(1)=6

MP(2)=17-6=11

MP(3)=27-17=10

MP(4)=32-27=5

MP(5)=30-32=-2


The diminishing marginal product set in when the number of workers is 3